# A BERTology View of LLM Orchestrations: Token- and Layer-Selective Probes for Efficient Single-Pass Classification

### v1

@@ A BERTology View of LLM Orchestrations:
Gonzalo Ariel Meyoyan Departamento de Computación, FCEyN Universidad de Buenos Aires Buenos Aires, Argentina gmeyoyan@dc.uba.ar          Luciano Del Corro ELIAS Lab, Departamento de Ingeniería Universidad de San Andrés Victoria, Argentina delcorrol@udesa.edu.ar
@@ 1 Introduction
This shifts the question from whether we can reuse the serving model’s computation to where in the computation we should read out: which tokens and which layers provide the most discriminative features for a given classification objective? Motivated by this view, we treat moderation and NLU classification as representation-selection problems over the full $L\times T\times d$ hidden-state tensor, and learn lightweight, aggregators that select and combine information across both tokens and layers while keeping the serving LLM frozen.
@@
We evaluate our probes on safety moderation and sentiment classification, using a relatively small Llama-3.2-3B-Instruct as the serving backbone. Across ToxicChat and WildGuardMix, our probes match or surpass strong guard baselines while requiring orders of magnitude fewer trainable parameters, and they remain competitive on IMDB, SST-2 and Emotion. Beyond accuracy, we show that reusing hidden states yields a simple deployment path: classification can be performed alongside generation with a single model call, avoiding the latency and overhead of a separate guard model.
@@ 2 Related Work
Orchestration and Guard Models. Production LLM systems typically complement the serving model with auxiliary classifiers for content moderation, policy compliance, and safety (Inan et al., 2023; Han et al., 2024; Ghosh et al., 2024). These guard models improve safety but add latency, memory, and deployment complexity. An alternative is safety aligning the main model, but overdoing it risks degrading the model’s overall capabilities, an effect sometimes referred to as the alignment tax (Huang et al., 2025). Our approach provides additional safety without training big models and avoiding complex deployment pipelines, requiring only 35M additional trainable parameters versus several billion for a separate guard model, while reusing the forward pass already performed by the serving LLM.
Reusing Computation for Classification. Recent work extracts classification signals from the LLM. MULI (Hu et al., 2024) trains a sparse classifier on first-token logits; LPM (Chrabąszcz et al., 2025) applies distance-based rules to a vector. For safety probing, ShieldHead (Xuan et al., 2025) attaches a head to final-layer hidden states, OmniGuard (Verma et al., 2025) probes a single layer, and Zhou et al. (2024) analyze how safety signals emerge across depth. These methods reduce orchestration cost but commit to a fixed layer or token, underutilizing signals distributed across the model. In contrast, we introduce an approach to search for signals across layer and token hidden states.
Layer Selection and Multi-Layer Aggregation. Interpretability work shows that transformer layers encode different abstractions: lower layers capture syntactic patterns while higher layers encode semantics (Tenney et al., 2019; Jawahar et al., 2019). Subsequent analyses suggest that task-relevant information is distributed across depth and that combining layers outperforms single-layer selection (de Vries et al., 2020; Rogers et al., 2020). There must be for each task, a common ancestor representation that is useful for both. This motivates our design choice: rather than commit to a fixed readout position, we learn aggregators that discover which layers and tokens are most informative for a given task. We propose a principled, data-driven way to select and combine representations for a classifier by learning lightweight aggregators over tokens and layers, rather than committing to a fixed depth or pooling strategy.
@@
Figure 1: View of our probing architecture. The aggregation method featured here is the scoring attention gate
@@ 3 Probe Architecture
Formally, given a frozen decoder-only LLM with $L$ layers and an input prompt $x$ tokenized into $T$ tokens, the model produces hidden states $\mathbf{h}^{(l)}\in\mathbb{R}^{T\times d}$ at each layer $l\in\{0,\dots,{L-1}\}$, where $d$ is the hidden dimension (with l=0l{=}0 denoting the embedding output). Our goal is to learn a lightweight classifier $C_{\theta}$ that predicts a label $y$ (e.g., safe/unsafe, positive/negative) from these representations, i.e., $y=C_{\theta}(\{\mathbf{h}^{(l)}\}_{l=0}^{L-1})$. The LLM remains frozen and we train only $\theta$, enabling retrofitting to deployed models while reusing the serving forward pass. The main challenge is to aggregate information across both tokens ($T$) and layers ($L$) to extract discriminative features for classification.
@@ 3.1 Two-Stage Aggregation Architecture
Our architecture addresses this dimensionality challenge with a two-stage aggregation scheme. The LLM produces a three-dimensional hidden-state tensor of shape $L\times T\times d$ (layers $\times$ tokens $\times$ hidden dimension). To avoid feeding this tensor directly to a lightweight classifier, we first aggregate across tokens within each layer to obtain $L$ fixed-size layer summaries, then aggregate across layers to produce a single $d$-dimensional vector that is consumed by a standard classification head.
@@
Figure 1 summarizes our pipeline. Given hidden states $\{\mathbf{h}^{(l)}\}_{l=0}^{L-1}$, we (Stage 1) aggregate tokens within each layer to obtain layer summaries $\mathbf{v}^{(l)}\in\mathbb{R}^{d}$, (Stage 2) aggregate across layers to produce a single vector $\mathbf{v}\in\mathbb{R}^{d}$, and then apply a linear head to obtain class logits. We use the same aggregation mechanism for both stages (Section 3.2).
@@
trained with cross-entropy.
@@ 3.2 Aggregation Mechanisms
Let $\mathbf{X}\in\mathbb{R}^{N\times d}$ denote the input to an aggregation block, where $N\in\{T,L\}$ depends on the stage. The block outputs a fixed-size vector $\mathbf{v}\in\mathbb{R}^{d}$ summarizing the most informative content in $\mathbf{X}$.
@@
For token-level aggregation, we use one gate per layer ($L$ total); for layer-level aggregation, we use a single shared gate. Parameter count: (L+1)d(L{+}1)d trainable parameters.
@@
As with the scoring gate, we use L+1L{+}1 MHA modules (one per layer for Stage 1, one for Stage 2), for a total of $(L{+}1)\cdot 4d\,d_{\text{inner}}$ parameters.
@@ 4 Experiments and Results
Table 3: WildGuardMix. Reuse methods attach a probe to a serving Llama-3.2-3B model and reuse the forward pass; we report added parameters beyond the backbone. Standalone baselines require an model invocation.
@@ 4.1 Experimental Setup
Model. We use Llama-3.2-3B-Instruct (Grattafiori et al., 2024) to simulate a serving LLM. [^1] Code available at https://anonymous.4open.science/r/LLM-Probing-Classifiers-2526 Notably, this compact 3B backbone is substantially smaller than many guard-model and classifier baselines, yet it already exposes hidden representations that support strong moderation and NLU probes on our benchmarks. In many real deployments, the serving model is substantially larger, and established scaling laws suggest that language modeling and downstream capability tend to improve with model size and data/compute (Kaplan et al., 2020; Hoffmann et al., 2022). Since our probes read out from the serving model’s internal representations without modifying the backbone, stronger backbones offer a natural path to improved probe accuracy at essentially unchanged relative probe overhead; we leave a systematic backbone sweep to future work. We focus on the Llama family for comparability and practical relevance, as it underlies widely used open guard-model baselines and is common in recent safety evaluations. We treat the scaling argument as a hypothesis rather than a guarantee, and empirically validating cross-backbone transfer is an important next step.
@@
WildGuardMix. WildGuardMix (Han et al., 2024) is a large safety benchmark (8̃9K examples) mixing multiple sources and synthetic data across diverse harm categories (e.g., violence, hate, self-harm, sexual content, privacy), with both direct requests and adversarial jailbreaks. We use the standard train/test split.
IMDB / SST-2 / Emotion. IMDB (Maas et al., 2011) is binary sentiment over movie reviews; following Palma et al. (2025), we use their 12,500-example subset (6,250/6,250). SST-2 (Socher et al., 2013) is binary sentiment over short snippets (standard split). Emotion (Saravia et al., 2018) is a multi-class emotion classification benchmark; following Palma et al. (2025), we use their 8,000-example subset.
@@ 4.3 Safety Classification Results
We evaluate whether safety signals in a frozen serving LLM can be read out with lightweight probes that run in the same forward pass as generation. To make the deployment trade-off explicit, we group baselines by whether they require an extra LM call at inference (standalone guard models / APIs) versus reuse methods that attach a small head to the serving model (MULI and ours). For all reuse methods, we report added parameters beyond the frozen Llama-3.2-3B-Instruct backbone.
ToxicChat: in-distribution (Table 1). When trained and evaluated on ToxicChat, learned token–layer aggregation improves over fixed readouts. Direct pooling reaches 73.53 F1, while the scoring-attention gate improves to 80.49 F1 with 0.10M added parameters. The multi-head self-attention (MHA) probe achieves 84.51 F1 and 0.898 AUPRC, improving over the logit-reuse baseline MULI (77.8 F1 / 0.829 AUPRC) and the standalone ToxicChat-T5 classifier (82.2 F1 / 0.885 AUPRC), while avoiding an extra model invocation. These gains are obtained with aggressive attention downcasting, suggesting that the improvement is driven primarily by where we read out (across tokens and layers) rather than by large additional capacity.
@@
Takeaway. Across both datasets, method ranking is consistent (pooling << scoring-attention << MHA). Overall, the benefit is not only reuse of the serving computation, but learning where in the $L\times T\times d$ tensor the safety signal is most separable, reducing the latency and VRAM overhead associated with a second LLM in the orchestration pipeline.
Figure 2: Attention weights from the scoring-attention probe on ToxicChat, stratified by label and correctness. The dashed line marks uniform $1/L$ ($\approx 0.034$). Toxic prompts attend to later layers (L17–L28), while non-toxic prompts concentrate on layers L0 and L27–L28. Misclassifications resemble the predicted-class profile.
@@ 6 Inference Efficiency Analysis
[…]
Deployment scenarios. Our baseline is a guard-then-serve pipeline in which Llama-Guard-3-8B (Inan et al., 2023) screens prompts and the model (Llama-3.2-3B-Instruct) is invoked after a guard decision. In contrast, our probes are attached to the model and compute the label from hidden states produced in the forward pass as generation, eliminating an extra model invocation.
Results. Table 5 shows a clear separation between single-pass reuse and a two-model pipeline. Relative to the serving-only baseline, pooling adds negligible overhead and the scoring-attention gate remains close. The MHA probe is slower, reflecting the additional attention computation, but still substantially faster than the guard-then-serve pipeline. Throughput follows the same trend.
Peak memory is dominated by whether a second model is loaded. All probe variants remain near the serving footprint (6.5–7.0 GB), whereas the guard-then-serve pipeline requires 22.8 GB due to the additional 8B guard model. Overall, these measurements show that (i) reuse-based classification preserves a single-model serving profile, and (ii) within our probe family, pooling and scoring-attention offer the best latency–accuracy trade-off, while MHA provides higher accuracy at a moderate but still single-pass overhead.
@@ 7 Conclusions and Future Work
We showed that lightweight probes trained on a frozen serving LLM’s hidden states can support moderation and sentiment/emotion classification without deploying a separate classifier model. By casting classification as representation selection over the full $L\times T\times d$ hidden-state tensor and using a two-stage token–layer aggregation, our approach avoids committing to a single layer or token while remaining parameter-efficient. Across ToxicChat and WildGuardMix, our probes approach or match strong guard-model baselines with far fewer trainable parameters, and they are competitive on IMDB, SST-2, and Emotion. System-wise, reuse enables a single-pass pipeline that reduces latency, VRAM, and orchestration complexity relative to guard-then-serve deployments. A key next step is to validate robustness across backbones.
@@
We evaluated our approach exclusively on Llama-3.2-3B-Instruct. While this demonstrates that even compact models encode rich hidden representations, generalization to other model families (e.g., GPT, Gemma, Mistral) and sizes (1B to 70B+) remains unexplored. Different architectures may encode task-relevant information differently across layers, potentially requiring architecture-specific tuning of our aggregation mechanisms. Furthermore, our approach inherits the semantic understanding, capabilities, and biases of the underlying frozen LLM. Since we train probes on the model’s hidden representations without modifying the base model, any biases present in the LLM’s pretraining data or limitations in its semantic understanding will directly affect classification performance.
@@
Unlike safety fine-tuning or guard models that can generate explanatory rejections (e.g., “I cannot help with that request because…”), our classification approach can only detect harmful content and interrupt generation. While this may be addressable through conditional re-prompting after detection, we did not evaluate such strategies. This limits the user experience compared to models that provide contextual explanations during content moderation.
@@
Understanding where safety signals are encoded in LLM hidden states could theoretically inform adversarial attacks designed to evade detection. However, similar information is available in prior interpretability work, and we believe the benefits of efficient moderation outweigh this risk.

### v2

@@ A BERTology View of LLM Orchestrations:
Gonzalo Ariel Meyoyan Departamento de Computación, FCEyN Universidad de Buenos Aires Buenos Aires, Argentina gmeyoyan@dc.uba.ar          Luciano Del Corro ELIAS Lab, Departamento de Ingeniería Universidad de San Andres Victoria, Argentina delcorrol@udesa.edu.ar
@@ 1 Introduction
This shifts the question from whether we can reuse the serving model’s computation to where in the computation we should read out: which tokens and which layers provide the most discriminative features for a given classification objective? Motivated by this view, we treat moderation and NLU classification as representation-selection problems over the full $L\times T\times d$ hidden-state tensor, and learn lightweight aggregators that select and combine information across both tokens and layers while keeping the serving LLM frozen.
@@
We evaluate our probes on safety moderation and sentiment classification using Llama-3.2-3B-Instruct as the primary serving backbone, and validate generalization on two additional architectures (GPT-OSS-20B and Qwen3-30B-A3B), spanning dense and mixture-of-experts designs. Across ToxicChat and WildGuardMix, our probes match or surpass strong guard baselines while requiring orders of magnitude fewer trainable parameters, and they remain competitive on IMDB, SST-2 and Emotion. The method ranking (pooling < scoring attention < MHA) is consistent across all three backbones. Beyond accuracy, we show that reusing hidden states yields a simple deployment path: classification can be performed alongside generation with a single model call, avoiding the latency and overhead of a separate guard model.
@@ 2 Related Work
Orchestration and Guard Models. Production LLM systems typically complement the serving model with auxiliary classifiers for content moderation, policy compliance, and safety (Inan et al., 2023; Han et al., 2024; Ghosh et al., 2024). These guard models improve safety but add latency, memory, and deployment complexity. An alternative is safety aligning the main model, but overdoing it risks degrading the model’s overall capabilities, an effect sometimes referred to as the alignment tax (Huang et al., 2025). Our approach provides additional safety without training big models or deploying complex pipelines, requiring only 35M additional trainable parameters versus several billion for a separate guard model, while reusing the forward pass already performed by the serving LLM.
Reusing Computation for Classification. Recent work extracts classification signals from the LLM. MULI (Hu et al., 2024) trains a sparse classifier on first-token logits; LPM (Chrabąszcz et al., 2025) applies distance-based rules to a fixed latent representation. For safety probing, ShieldHead (Xuan et al., 2025) attaches a head to final-layer hidden states, OmniGuard (Verma et al., 2025) probes a single layer, and Zhou et al. (2024) analyze how safety signals emerge across depth. These methods reduce orchestration cost but commit to a fixed layer or token, underutilizing signals distributed across the model. In contrast, we introduce an approach to search for signals across layer and token hidden states.
Layer Selection and Multi-Layer Aggregation. Interpretability work shows that transformer layers encode different abstractions: lower layers capture syntactic patterns while higher layers encode semantics (Tenney et al., 2019; Jawahar et al., 2019). Subsequent analyses suggest that task-relevant information is distributed across depth and that combining layers outperforms single-layer selection (de Vries et al., 2020; Rogers et al., 2020). This premise has also been validated outside of classification: in speculative decoding, EAGLE-3 (Li et al., 2025) shows that fusing features from multiple transformer depths substantially outperforms top-layer-only representations for draft-token prediction. These findings motivate our design choice: rather than commit to a fixed readout position, we learn aggregators that discover which layers and tokens are most informative for a given task. We propose a principled, data-driven way to select and combine representations for a classifier by learning lightweight aggregators over tokens and layers, rather than committing to a fixed depth or pooling strategy.
@@
Figure 1: Overview of the two-stage probing architecture, illustrated with the scoring attention gate. Stage 1 summarizes tokens within each layer; Stage 2 aggregates across layers to produce a single classification vector.
@@ 3 Probe Architecture
Formally, given a frozen decoder-only LLM with $L$ layers and an input prompt $x$ tokenized into $T$ tokens, the model produces hidden states $\mathbf{h}^{(l)}\in\mathbb{R}^{T\times d}$ at each layer $l\in\{0,\dots,{L}\}$, where $d$ is the hidden dimension (with l=0l{=}0 denoting the embedding output). Our goal is to learn a lightweight classifier $C_{\theta}$ that predicts a label $y$ (e.g., safe/unsafe, positive/negative) from these representations, i.e., $y=C_{\theta}(\{\mathbf{h}^{(l)}\}_{l=0}^{L})$. The LLM remains frozen and we train only $\theta$, enabling retrofitting to deployed models while reusing the serving forward pass. The main challenge is to aggregate information across both tokens ($T$) and layers ($L$) to extract discriminative features for classification.
@@ 3.1 Two-Stage Aggregation Architecture
Our architecture addresses this dimensionality challenge with a two-stage aggregation scheme. The LLM produces a three-dimensional hidden-state tensor of shape $L\times T\times d$ (layers $\times$ tokens $\times$ hidden dimension). To avoid feeding this tensor directly to a lightweight classifier, we first aggregate across tokens within each layer to obtain $L+1$ fixed-size layer summaries, then aggregate across layers to produce a single $d$-dimensional vector that is consumed by a standard classification head.
@@
Figure 1 summarizes our pipeline. Given hidden states $\{\mathbf{h}^{(l)}\}_{l=0}^{L}$, we (Stage 1) aggregate tokens within each layer to obtain layer summaries $\mathbf{v}^{(l)}\in\mathbb{R}^{d}$, (Stage 2) aggregate across layers to produce a single vector $\mathbf{v}\in\mathbb{R}^{d}$, and then apply a linear head to obtain class logits. We use the same aggregation mechanism for both stages (Section 3.2).
@@
trained with cross-entropy. At inference time, the probe emits a label from the same forward pass that initiates generation. If the input is classified as unsafe, the orchestration layer can halt generation before any tokens are streamed and either return a templated refusal or re-prompt the serving LLM to produce a contextual rejection, with no additional model invocation.
@@ 3.2 Aggregation Mechanisms
Let $\mathbf{X}\in\mathbb{R}^{N\times d}$ denote the input to an aggregation block, where $N\in\{T,L+1\}$ depends on the stage. The block outputs a fixed-size vector $\mathbf{v}\in\mathbb{R}^{d}$ summarizing the most informative content in $\mathbf{X}$.
@@
For token-level aggregation, we use one gate per layer ($L+1$ total); for layer-level aggregation, we use a single shared gate. Parameter count: (L+2)d(L{+}2)d trainable parameters.
@@
As with the scoring gate, we use L+2L{+}2 MHA modules (one per layer for Stage 1, one for Stage 2), for a total of $(L{+}2)\cdot 4d\,d_{\text{inner}}$ parameters.
@@ 4 Experiments and Results
GPT-OSS-20B backbone
   Direct pooling 77.36 $\pm$ 0.87 0.833 $\pm$ 0.008 0.003
   Scoring attention 79.23 $\pm$ 0.58 0.854 $\pm$ 0.005 0.08
   Multi-head self-attn 86.17 $\pm$ 0.51 0.915 $\pm$ 0.004 27
Qwen3-30B-A3B backbone
   Direct pooling 73.24 $\pm$ 0.52 0.827 $\pm$ 0.005 0.002
   Scoring attention 80.94 $\pm$ 0.62 0.883 $\pm$ 0.006 0.10
   Multi-head self-attn 83.76 $\pm$ 0.9 0.905 $\pm$ 0.005 26
@@
Table 3: WildGuardMix. Reuse methods attach a probe to a serving Llama-3.2-3B model and reuse the forward pass; we report added parameters beyond the backbone. Standalone baselines require an additional model invocation.
@@ 4.1 Experimental Setup
[…]
@@
WildGuardMix. WildGuardMix (Han et al., 2024) is a large safety benchmark ($\sim$ 89K examples) mixing multiple sources and synthetic data across diverse harm categories (e.g., violence, hate, self-harm, sexual content, privacy), with both direct requests and adversarial jailbreaks. We use the standard train/test split.
IMDB / SST-2 / Emotion. IMDB (Maas et al., 2011) is binary sentiment over movie reviews; following Palma et al. (2025), we use their 12,500-example subset (6,250/6,250). SST-2 (Socher et al., 2013) is binary sentiment over short snippets; following Palma et al. (2025), we use their 7,000-example subset. Emotion (Saravia et al., 2018) is a multi-class emotion classification benchmark; following Palma et al. (2025), we use their 8,000-example subset.
@@ 4.3 Safety Classification Results
We evaluate whether safety signals in a frozen serving LLM can be read out with lightweight probes that run in the same forward pass as generation. To make the deployment trade-off explicit, we group baselines by whether they require an extra LM call at inference (standalone guard models / APIs) versus reuse methods that attach a small head to the serving model (MULI and ours). For all reuse methods, we report added parameters beyond those from the frozen backbone.
ToxicChat: in-distribution (Table 1). When trained on the ToxicChat training set and evaluated on the held-out test set (restricted to human-annotated samples), learned token–layer aggregation improves over fixed readouts. Direct pooling reaches 73.53 F1, while the scoring-attention gate improves to 80.49 F1 with 0.10M added parameters. The multi-head self-attention (MHA) probe achieves 84.51 F1 and 0.898 AUPRC, improving over the logit-reuse baseline MULI (77.8 F1 / 0.829 AUPRC) and the standalone ToxicChat-T5 classifier (82.2 F1 / 0.885 AUPRC), while avoiding an extra model invocation. These gains are obtained with aggressive attention downcasting, suggesting that the improvement is driven primarily by where we read out (across tokens and layers) rather than by large additional capacity.
Cross-backbone consistency (Table 1). To test whether these findings depend on a specific model family, we repeat the ToxicChat evaluation on two mixture-of-experts backbones: GPT-OSS-20B and Qwen3-30B-A3B. The method ranking (pooling << scoring-attention << MHA) is preserved across all three architectures. All MHA probes match or exceed the standalone ToxicChat-T5-large baseline (82.2 F1) without an extra model call, with GPT-OSS-20B achieving the highest F1 (86.17) and AUPRC (0.915). These results suggest that the layer-distributed safety signal is neither an artifact of the Llama architecture nor of dense transformer structure, as MoE models (which route tokens through sparse expert subsets) exhibit the same pattern.
@@
Takeaway. Across both safety datasets, method ranking is consistent (pooling << scoring-attention << MHA), and multi-backbone experiments on ToxicChat confirm this ordering generalizes across dense and MoE architectures. Overall, the benefit is not only reuse of the serving computation, but learning where in the $L\times T\times d$ tensor the safety signal is most separable, reducing the latency and VRAM overhead associated with a second LLM in the orchestration pipeline.
Figure 2: Attention weights from the scoring-attention probe on ToxicChat, stratified by label and correctness. The dashed line marks uniform $1/(L+1)$ ($\approx 0.034$). Toxic prompts attend to later layers (L17–L28), while non-toxic prompts concentrate on layers L0 and L27–L28. Misclassifications resemble the predicted-class profile.
@@ 6 Inference Efficiency Analysis
+ ToxicChat-T5 (780M) 11.32 88.33 7992
+ WildGuard (7B) 9.97 100.32 21084
+ Llama Guard 3 (8B) 8.12 123.21 22769
Table 5: Latency benchmark on ToxicChat test prompts using Llama-3.2-3B-Instruct as the serving model (2000 samples, max length 512, batch size 1, max_new_tokens=1). Probe rows reuse the serving forward pass. Guard-model rows (ToxicChat-T5, WildGuard, Llama Guard 3) run a separate model followed by the serving LLM, covering the architectural families behind the majority of open-source guard models presented in our previous safety benchmarks.
Deployment scenarios. Our baseline is a guard-then-serve pipeline in which a separate guard model screens prompts before the serving Llama-3.2-3B-Instruct is invoked. We benchmark three representative guard models: ToxicChat-T5 (780M), WildGuard (7B, Mistral fine-tune), and Llama Guard 3 (8B) (Grattafiori et al., 2024). In contrast, our probes are attached to the serving model and compute the label from hidden states produced during the same forward pass, eliminating an extra model invocation.
@@
Peak memory is dominated by whether a second model is loaded. All probe variants remain near the serving footprint (6.5–7.0 GB), whereas guard-then-serve pipelines range from 8.0 GB (ToxicChat-T5) to 22.8 GB (Llama Guard 3) depending on guard model size. Overall, these measurements show that (i) reuse-based classification preserves a single-model serving profile, and (ii) within our probe family, pooling and scoring-attention offer the best latency–accuracy trade-off, while MHA provides higher accuracy at a moderate but still single-pass overhead.
@@ 7 Conclusions and Future Work
We showed that lightweight probes trained on a frozen serving LLM’s hidden states can support moderation and sentiment/emotion classification without deploying a separate classifier model. By casting classification as representation selection over the full $L\times T\times d$ hidden-state tensor and using a two-stage token–layer aggregation, our approach avoids committing to a single layer or token while remaining parameter-efficient. Across ToxicChat and WildGuardMix, our probes approach or match strong guard-model baselines with far fewer trainable parameters, and they are competitive on IMDB, SST-2, and Emotion. Multi-backbone experiments on dense and mixture-of-experts architectures confirm that the method ranking and competitive performance are not specific to a single model family. System-wise, reuse enables a single-pass pipeline that reduces latency, VRAM, and orchestration complexity relative to guard-then-serve deployments. Key next steps include extending multi-backbone evaluation to additional tasks, testing with larger active parameter counts, and exploring fine-grained safety taxonomies.
@@
We validated cross-architecture robustness on ToxicChat using three backbone families (Llama-3.2-3B, GPT-OSS-20B, Qwen3-30B-A3B), but sentiment and emotion benchmarks were evaluated only on Llama-3.2-3B-Instruct. Moreover, all three backbones share approximately 3B active parameters, so the effect of substantially larger active capacity on probe accuracy remains untested. Different architectures may encode task-relevant information differently across layers, potentially requiring architecture-specific tuning of our aggregation mechanisms.
@@
Unlike safety fine-tuning or guard models that can generate explanatory rejections (e.g., “I cannot help with that request because…”), our classification approach can only detect harmful content and interrupt generation. However, because our probes complement rather than replace the serving LLM, the orchestration layer can trigger a refusal response from the same model once unsafe content is detected; the serving model’s generation capability is available by design. We did not evaluate such conditional re-prompting strategies, and integrating probe-based detection with contextual refusal generation is a natural direction for future work.
Robustness and Failure Scenarios.
Our OOD results (Table 2) show that distribution shift degrades probe performance. Because probes are lightweight classifiers over learned feature patterns, they may fail silently on novel harm categories, evolving jailbreak strategies, or input distributions not represented in training data. Unlike full guard LLMs, which can reason over text and generalize more broadly, probes cannot adapt without retraining. In safety-critical deployments with rapidly evolving threat landscapes, or in domains requiring coverage of harm categories beyond the training distribution, a dedicated guard model may remain necessary despite its higher computational cost.
@@
Understanding where safety signals are encoded in LLM hidden states could inform adversarial attacks designed to evade detection. However, similar information is available in prior interpretability work, and we believe the benefits of efficient moderation outweigh this risk.